# Transport through a lattice with local loss: from quantum dots to lattice gases

### v1

@@ I.1 Motivation and background
Reference [2] has solved this general problem with nonequilibrium Green’s functions in the Keldysh formalism extended to open quantum systems. The main result for the steady-state fermionic current through the finite chain is that conductivity shows discrete steps, as a function of voltage between the reservoirs across the chain of sites, and that the lossy site systematically smooths out these steps, even with moderate loss rates. Higher loss rates also significantly lower the steps, reducing overall conductivity.
Letting the single lossy site be the central site, with an odd number of sites in the chain, elegantly isolates the effect of loss on the conductivity steps. The conductivity steps are associated with energetic access to single-particle eigenstates in the finite chain; with reflection symmetry, odd-parity eigenstates have nodes at the loss site, and transport through these channels is much less affected by the loss than transport through even-parity channels. This shows up clearly in [2] in conductivity steps that alternate between smooth and steep.
Although [2] explicitly raises the possibility of experimental realizations with ultracold fermi gases in optical lattices, however, its concrete results are obtained for parameter regimes that rather represent electron transport through a chain of quantum dots, with the reservoir roles played by macroscopic wire leads. In this paper we therefore look critically at a quantum gas realization of this lossy transport problem. Instead of applying an open-system formalism, we replace the macroscopic fermion reservoirs of [2] with more lattice sites like the ones in the finite middle chain. The finite middle chain is only separated from the “reservoir” parts of the lattice by “weak links” of slower tunneling rates between sites. We will show, first of all, that this particular class of Hamiltonian realizations of the particle reservoirs can indeed reproduce exactly the results of [2]. We will then re-tune our parameters to represent experiments that will likely be more straightforward with quantum gases. We will investigate both finite-temperature and finite-size effects, and conclude that experiments with quantum gases are realistically feasible, and can provide complementary insights to those offered by quantum dots.
@@ I.2 Paper structure
In Section IV we will then re-tune our infinite model to correspond more closely to the most straightforward kind of quantum gas experiments, with chemical potential differences between the left and right reservoirs, but no on-site potential differences, and with densities of states in the middle and outer sections of the lattice all equal, instead of having much greater densities of states in the reservoirs. We predict steady-state currents that are qualitatively similar to those found in [2] for the quantum dot system with leads. There are abrupt steps in the conductivity at chemical potential differences corresponding to eigenfrequencies of the isolated middle lattice; the steps are smoothed by the loss at the central site, or remain quite sharp, depending on the parity of the corresponding eigenfunctions. Although qualitatively very similar, the straightforward quantum gas scenario is found to differ in detail from the quantum dot version of the experiment.
@@ II.1 Infinite lattices with weak links as finite lattices between reservoirs
The site-dependent parameters $T_{m}>0$ and $V_{m}$ do not depend arbitrarily on $m$, but have a simple pattern that defines a symmetrical finite chain of $12M+$ 1 sites ($|m|\leq M$) centered between semi-infinite reservoir zones on the left and right:
@@ II.3.1 Single-particle observables
$\sum_{m}R_{m^{\prime}m}=\mathrm{Tr}\Big{(}\hat{\rho}\sum_{m}\hat{a}^{\dagger}_{m^{\prime}}\hat{a}_{m}\Big{)}\;.$ (12)
@@ III.2 Left- and right-incident normalized eigenfunctions
$\sqrt{2\pi}\Psi_{n}^{L}(\omega)$ =$=\left\{\begin{array}[]{ccc}e^{ik_{L}m}+\mathcal{R}_{L}(\omega)e^{-ik_{L}m}&\;,&m<-M\\ \sqrt{\frac{sin(k_{L})}{sin(k_{R})}}\mathcal{T}(\omega)e^{ik_{R}m}&\;,&m>M\end{array}\right.$ (19)
$\sqrt{2\pi}\Psi_{n}^{R}(\omega)$ =$=\left\{\begin{array}[]{ccc}\sqrt{\frac{sin(k_{R})}{sin(k_{L})}}\mathcal{T}(\omega)e^{-ik_{L}m}&\;,&m<-M\\ e^{-ik_{R}m}+\mathcal{R}_{R}(\omega)e^{ik_{R}m}&\;,&m>M\end{array}\right.\;,$ (22)
@@ III.3 Why scattering wave functions?
The first reason for defining the left- and right-incident wave functions $\Psi_{m}^{L,R}$ in this way (19) is that they allow us easily to create arbitrary incident wave packets at infinity. For example, in a superposition of the form
[…]
for arbitrary $\bar{k}\in[0,\pi]$ and large $\Delta m\gg 1$ but even larger negative $\bar{m}\ll-\Delta m$, only the incoming wave term $\propto e^{ik_{L}m}$ in $\Psi_{m}^{L}$ will contribute significantly—and it will contribute essentially the same Gaussian wave packet, centered on $m=\bar{m}$ and wave number $k_{L}=\bar{k}$, that we would obtain if $\Psi_{m}^{L}$ consisted of nothing but that incoming wave because we were dealing with an infinite uniform lattice with no weak or lossy links anywhere.
It is then the implication of these incoming wave packets from infinity that is our main reason for defining the $\Psi_{m}^{L,R}$. An equilibrium ensemble of incoming wave packets represents the particles which approach our finite middle chain, with its lossy site, from an infinite reservoir of particles in equilibrium. Consequently we must be able to represent the single particle density matrix $R_{m^{\prime}m}$ of a steady state of our system, with distinct equilibrium distributions at left and right infinity, in the form (16) that we defined at the beginning of this Section.
@@ III.6.1 Transmission and reflection coefficients $\mathcal{T}$ and $\mathcal{R}$ for $\Lambda\to\infty$
For $\Lambda\to\infty$ the continuous bandwidth $|\omega\mp V/2|\leq 2\Lambda$ of the semi-infinite outer parts of the lattice becomes much broader than the bandwidth $|\omega|\leq 2$ of the finite middle chain of sites. Wave vectors for $|\omega|\gg 2$ decay so sharply within $|m|<M$ that they contribute negligibly to any observables within the middle chain, including the particle current through the middle chain, so as far as any such observables are concerned, we can restrict our attention to $|\omega|\ll\Lambda$. For all these $\omega$, the limit $\Lambda\to\infty$ means that $k_{L,R}$ are both negligibly different from $\pi/2$. This implies that the density of states in the “reservoir” ends of the lattice is uniform over the entire frequency range that matters for transport through the finite middle chain.
@@ IV.1 The quantum gas set-up
This less symmetrical but simpler set-up also has the advantage of showing more conductivity steps. In the ±$\pm V\not=0$ scenario with equal fermion densities in leads, the currents from right and from left exactly cancel in the frequency range $\omega<-V/2$, so that the integral (29) over $\omega$ is effectively over the symmetrical middle range $|\omega|<V/2$. Since the resonant frequencies of the transmission channels through the finite middle chain are in symmetrical pairs ±$\pm\omega_{n}$, raising $V$ effectively only opens new channels in pairs, producing only $M+1$ conductivity steps. In the quantum gas scenario with particles coming only from the left reservoir, in contrast, raising $\mu_{L}$ through the system’s frequency band sweeps through each of the $12M+$ 1 transmission resonances separately.
Figure 2: Dimensionless fermion current $I$ at zero temperature, as a function of left-side chemical potential $\mu_{L}$, for two different central-site loss rates $\gamma$. Both plots show $\tau_{1}=0.1$ as in Fig. 1, but with $V=0$, $\Lambda=1$, and no fermions incident from the right. Although the middle chains here have seven sites ($M=3$), the leftmost and rightmost current steps are small and gentle in a), and completely flattened in b), leaving only five current steps clearly visible. In a) we have $\gamma=0.5$, while b) shows $\gamma=5$. As well as suppressing the outermost of the seven current steps, the moderate loss in a) has noticeably smoothed out two of the five more visible steps in the current, while the stronger loss in b) has smoothed them completely into gradual slopes. The other three visible current steps are not significantly affected by loss in either case, since they are due to transport through odd-parity channels having nodes at the lossy site $m=0$. Although $I(\mu_{L})$ differs in detail from analogous plots of $I(V)$ for quantum dots between leads, the same phenomena of conductivity steps with parity-dependent smoothing by loss are revealed.
@@
Five steplike rises in conductivity through the finite middle chain of seven sites ($M=3$) are clearly visible in Fig. 2; looking closely at a), one can also discern two more small, smooth steps at the left and right edges of the plot. Also apparent is the effect of loss in making some steps noticeably steeper than others. In Fig. 2a) the tiny first and seventh steps, as well as the more visible third and fifth steps, are all due to even-parity channels that are affected by the loss at $m=0$, and so more gently sloped than the three other steeper steps, which are due to odd-parity channels with nodes at $m=0$. The ten-times-higher loss rate in Fig. 2b) is enough to suppress the first and seventh steps entirely, and eliminate the shoulders of what were the third and fifth steps in a), leaving them as gradual slopes, while still leaving the three parity-protected transport resonances as sharp steps.
@@ IV.2 Finite temperature effects
Since the main phenomena to be studied in this kind of transport are the conductivity steps and their parity- and $\gamma$-dependent steepness, thermal smearing of the conductivity curve is a serious issue for experiments. Since the frequency spacing of our transmission resonances is of order $2/(2M+1)$, temperatures higher than this spacing (in units of $\hbar\tau/k_{B}$) will wash out the conductivity steps. Even if $T/\tau$ is below the $2/(2M+1)$ threshold, moreover, thermal smoothing of the conductivity steps can reduce their dramatic signature of discrete transmission channels to only a slight waviness in $I(\mu_{L})$, as seen in Fig. 3a); in this regime, moreover, the additional smoothing of even-parity channels due to the $m=0$ loss can be hard to discern. Both the steps and the effect of parity can still be seen at sufficiently low temperatures, as in Fig. 3b), but such low temperatures may be experimentally challenging. If the basic $T<2\tau/(2M+1)$ limit can be beaten, the problem of losing visibility of the parity effect for $T\lesssim\gamma$ can be solved by increasing $\gamma$, as in Fig. 3c).
@@ V.2 Quasi-stationary currents
For example, a natural initial state to prepare experimentally is a finite-system analog of the current-from-the-left scenarios shown for infinite lattices in Fig. 2. To prepare the initial state we set $\tau_{1}\to 0$, fully breaking the weak link and isolating the middle chain of our lattice from its outer ends. We then populate the isolated left end of the lattice with fermions in equilibrium, leaving the middle and right-end lattices empty. At $t=0$ we then restore non-zero $\tau_{1}$ and let the left-end population of fermions flow through the weakly linked middle chain, with its loss at $m=0$, and into the right lattice end. At different times $t>0$ we measure the decreasing population difference between left and right sides,
@@
As we can see in Fig. 4a), there is a brief initial period in which the fermions first begin to move through the middle chain of sites, via the weak links. Within a time of order $M/\tau$, however, this initial onset phase ends and we can see a remarkably steady fall of $\Delta\mathcal{N}$. The initial equilibrium state in the left end of the lattice can be viewed as an ensemble of wave packets moving in both directions; these packets move to and through the middle chain from further back in the left reservoir, just as they would if the reservoir were infinite. Even the packets which begin moving to the left, and then approach the middle chain after reflecting from the left end of the lattice, simply reproduce the effect of incoming packets from beyond the lattice end. The first effect which does not look just like an infinite system, in fact, is the return to the weak link of packets which have been reflected from it, and then reflected again from the left end of the lattice. This occurs for packets with group velocity $v_{G}$ at time $t=2N/v_{G}$. The maximum group velocity in our system is $2\tau$, occurring at $\omega=0$, and so the earliest finite size effects are those that appear in Fig. 4 at $t=N/\tau$. Between $t\lesssim M/\tau$ and $t<N/\tau$ there can be plenty of time to observe a quasi-stationary current, as long as $N$ is sufficiently larger than $M$. And as Fig. 4 shows, the rate at which $\Delta\mathcal{N}$ falls, during this quasi-stationary epoch between the initial transient and the onset of finite-size effects, can closely match the steady-state current of the infinite system.
@@
Figure 4: Finite-time evolution in finite systems. The initial state is prepared in a finite lattice with $N=20$, 40, or 80 sites in the outer lattice on each side. With the weak links between reservoir ends and middle chain initially turned off completely ($\tau_{1}\to 0$), the $N$-site left end lattice is initially half-filled at zero temperature, with the right end and central chain of seven sites ($M=3$) both initially empty. At $t=0$ the weak links are turned on $\tau_{1}\to 0.1$ and the system’s single-particle density matrix is evolved in time under (II.3.2) numerically, for a duration $2.5N/\tau$. The difference in total left- and right-side populations $\Delta N/2=(N_{L}-N_{R})/2$ is plotted in each case of $N$. The dashed line has the slope $-I$, where $I$ is the steady-state current for the $N\to\infty$ limit as shown in Fig. 2. The early transient epoch is so brief that it can only be seen clearly with shorter time axis of the $N=20$ plot. The onset of finite-size effects at $t=N/\tau$ is clear in all cases: with group velocity $2\tau$ for half filling, this is how long it takes for particles that initially reflect from the weak link to return to the weak link after reflecting again from the end of the lattice. For times after the short initial transient and before $t=N/\tau$, the finite system’s particle difference closely reproduces the infinite system’s steady-state current—in the case $\mu_{L}=0$ of initial half-filling. See Fig. 5 for other cases of $\mu_{L}$.

### v2

@@ I.1 Motivation and background
Reference [10] has solved this general problem with nonequilibrium Green’s functions in the Keldysh formalism extended to open quantum systems. The main result for the steady-state fermionic current through the finite chain is that conductance shows discrete steps, as a function of voltage between the reservoirs across the chain of sites, and that the lossy site systematically smooths out these steps, even with moderate loss rates. Higher loss rates also significantly lower the steps, reducing overall conductance.
Letting the single lossy site be the central site, with an odd number of sites in the chain, elegantly isolates the effect of loss on the conductance steps. The conductance steps are associated with energetic access to single-particle eigenstates in the finite chain; with reflection symmetry, odd-parity eigenstates have nodes at the loss site, and transport through these channels is much less affected by the loss than transport through even-parity channels. This shows up clearly in [10] in conductance steps that alternate between smooth and steep.
Although [10] explicitly raises the possibility of experimental realizations with ultracold fermi gases in optical lattices, however, its concrete results are obtained for parameter regimes that rather represent electron transport through a chain of quantum dots, with the reservoir roles played by macroscopic wire leads. Electron tunneling between quantum dots can certainly be realized [11], and one-dimensional chains of single-fermion sites provide idealized theoretical models for transport through quantum dot arrays[9, 12, 13]. In fact, realizing such ideally uniform chains of quantum dots is a longstanding technological challenge[14, 15, 16]. Although microscopic control of transport in solid state systems may have important applications in the future, the simplest cases of dissipative control over transport are currently more likely to be seen in lattice-trapped quantum gases.
In this paper we therefore look critically at a quantum gas realization of this lossy transport problem. Instead of applying an open-system formalism, we replace the macroscopic fermion reservoirs of [10] with more lattice sites like the ones in the finite middle chain; see Fig. 1. The finite middle chain is only separated from the “reservoir” parts of the lattice by “weak links” of slower tunneling rates between sites.
Figure 1: Spinless fermions move in a long chain of sites by nearest-neighbor tunneling; site-dependent tunneling rates define a finite chain of sites in the middle, separated by weak links from the long outer ends of the chain. Fermions can be lost from the central site. The system can model quantum dots between leads or cold atoms in an optical lattice.
We will show, first of all, that this particular class of Hamiltonian realizations of the particle reservoirs can indeed reproduce exactly the results of [10]. We will then re-tune our parameters to represent experiments that will likely be more straightforward with quantum gases. We will investigate both finite-temperature and finite-size effects, and conclude that experiments with quantum gases are realistically feasible, and can provide complementary insights to those offered by quantum dots.
@@ I.2 Paper structure
In Section IV we will then re-tune our infinite model to correspond more closely to the most straightforward kind of quantum gas experiments, with chemical potential differences between the left and right reservoirs, but no on-site potential differences, and with densities of states in the middle and outer sections of the lattice all equal, instead of having much greater densities of states in the reservoirs. We predict steady-state currents that are qualitatively similar to those found in [10] for the quantum dot system with leads. There are abrupt steps in the conductance at chemical potential differences corresponding to eigenfrequencies of the isolated middle lattice; the steps are smoothed by the loss at the central site, or remain quite sharp, depending on the parity of the corresponding eigenfunctions. Although qualitatively very similar, the straightforward quantum gas scenario is found to differ in detail from the quantum dot version of the experiment.
@@ II.1 Infinite lattices with weak links as finite lattices between reservoirs
The site-dependent parameters $T_{m}>0$ and $V_{m}$ do not depend arbitrarily on $m$, but have a simple pattern that defines a symmetrical finite chain of $12M+$ 1 sites ($|m|\leq M$) centered between semi-infinite reservoir zones on the left and right, as depicted in Fig. 1: we consider hopping rates and potentials of the form
@@ II.3.1 Single-particle observables
$\sum_{m}R_{mm}=\mathrm{Tr}\Big{(}\hat{\rho}\sum_{m}\hat{a}^{\dagger}_{m}\hat{a}_{m}\Big{)}\;.$ (6)
@@ II.3.3 Single-particle Schrödinger equation
In fact it can be shown that our Ansatz (II.3.3) includes the general solution to (II.3.2) as an initial value problem. See Appendix D for further discussion.
@@ III.2 Left- and right-incident normalized eigenfunctions
$\sqrt{2\pi}\Psi_{m}^{L}(\omega)$ =$=\left\{\begin{array}[]{ccc}e^{ik_{L}m}+\mathcal{R}_{L}(\omega)e^{-ik_{L}m}&\;,&m<-M\\ \sqrt{\frac{sin(k_{L})}{sin(k_{R})}}\mathcal{T}(\omega)e^{ik_{R}m}&\;,&m>M\end{array}\right.$ (13)
$\sqrt{2\pi}\Psi_{m}^{R}(\omega)$ =$=\left\{\begin{array}[]{ccc}\sqrt{\frac{sin(k_{R})}{sin(k_{L})}}\mathcal{T}(\omega)e^{-ik_{L}m}&\;,&m<-M\\ e^{-ik_{R}m}+\mathcal{R}_{R}(\omega)e^{ik_{R}m}&\;,&m>M\end{array}\right.\;,$ (16)
@@ III.3 Why scattering wave functions?
The reason for defining these left- and right-incident wave functions $\Psi_{m}^{L,R}$ in this way (13) is that they allow us to describe left and right reservoirs with different temperatures or chemical potentials, by considering separate grand canonical populations of the second-quantized $L$ and $R$ modes.
It may just seem obvious that uncorrelated populations of second-quantized $\Psi^{L}_{m}$ and $\Psi^{R}_{m}$ modes represent fermions approaching our middle chain from different equilibrium conditions. In fact the subject is worth some discussion. The set of all $\Psi_{m}^{L}$ and $\Psi_{m}^{R}$ for all $\omega$ together (plus any discrete bound states—see Appendix D) is complete, so any state can be represented in terms of them; but superpositions of $\Psi_{m}^{L}$ and $\Psi_{m}^{R}$ can also provide a complete set of states, and in general these superpositions will imply correlations and even coherences between incident particles from the left and incident particles from the right. It may be a physically natural condition to rule out such correlations, but it is still a non-trivial condition.
We can confirm in the following way that the kind of non-equilibrium steady state that we wish to consider in $R_{m^{\prime}m}$ is indeed the sum of a grand canonical distribution over $\psi_{m}^{L}$ modes and a second grand canonical distribution over $\psi_{m}^{R}$ modes. Suppose that we prepare the initial state of our infinite system by starting with $\tau_{1}\to 0$, so that all $\mathcal{T}(\omega)\to 0$ and all $|\mathcal{R}_{j}(\omega)|\to 1$. The left and right ends of the lattice are now completely separate semi-infinite uniform lattices, and the set of all $\Psi_{m}^{L}$ modes is a complete set of second-quantized normal modes in the left semi-infinite lattice, while the set of $\Psi_{m}^{R}$ modes is complete in the right.
We throw infinitely many particles into both of these separate semi-infinite lattices, by populating the $\Psi_{m}^{L}$ and $\Psi_{m}^{R}$ modes in second quantization. We then wait as long as it takes for the particles to reach equilibrium, either through contact with external baths which will later remove, or through interactions among the particles which are weak enough to have negligible effects on the nature of the steady state but which will thermalize the system over long enough times.
We now slowly raise $\tau_{1}$ to its desired steady-state value, so slowly that the $\Psi_{m}^{L,R}$ will all adiabatically evolve into their finite-$\tau_{1}$ forms, while keeping their average occupation numbers invariant. Over this long adiabatic ramping of $\tau_{1}$, whatever finite initial population was in the finite middle chain will either be lost at the impurity or else escape away to infinity through the weak links, while the infinite populations in the $\Psi_{m}^{L,R}$ so that the population remaining in finite $m$ sites is entirely given by the two grand canonical populations of $\Psi_{m}^{L,R}$, as assumed in (10).
Realistic state preparations may not take infinite time, or involve literally infinite systems, but they are quite likely to conform qualitatively to this adiabatic scenario. In any case, the steady state that results from the adiabatic $\tau_{1}$-ramping scenario must clearly be the kind of state that we mean when we speak of steady flow between reservoirs at two different energies and chemical potentials. If there is any other kind of state preparation that will produce such a state, the state itself will still be the same, up to the small errors that are always involved in treating large but finite systems in quasi-steady states as if they were literally in equilibrium.
@@ III.6.1 Transmission and reflection coefficients $\mathcal{T}$ and $\mathcal{R}$ for $\Lambda\to\infty$
Since the outer ends of the lattice are just large, uniform lattices, propagating modes with real $k_{L,R}$ can have frequencies within the continuous bands $|\omega\mp V/2|\leq 2\Lambda$, respectively for the left and right ends. For large $\Lambda$, these bands become very wide compared to the range of frequencies $|\omega|\leq 2$ within which the finite middle chain of sites can support real $k$ modes. Modes with $|\omega|\gg 2$ can propagate in the outer ends, but they must decay so sharply within $|m|<M$ that they contribute negligibly to any observables within the middle chain, including the particle current through the middle chain. As far as any such observables are concerned, therefore, we can certainly restrict our attention to $|\omega|\ll\Lambda$ when $\Lambda$ is large. For all these $|\omega|\ll\Lambda$, then, the limit $\Lambda\to\infty$ means that $k_{L,R}(\omega)$ are both negligibly different from $\pi/2$.
This lets us simplify our generally complicated formulas by replacing $k_{L,R}\to\pi/2$. The physical meaning of this mathematical convenience is that that for infinite $\Lambda$, the density of states $1/(dk_{L,R}/d\omega)$ in the “reservoir” ends of the lattice is uniform over the entire frequency range that matters for transport through the finite middle chain. This reproduces the uniform density of states that was assumed for the external leads in [10], and this can be considered the ultimate reason why the $\Lambda\to\infty$ limit lets our infinite lattice model duplicate the open system.
@@ IV.1 The quantum gas set-up
This less symmetrical but simpler set-up also has the advantage of showing more conductance steps. In the ±$\pm V\not=0$ scenario with equal fermion densities in leads, the currents from right and from left exactly cancel in the frequency range $\omega<-V/2$, so that the integral (22) over $\omega$ is effectively over the symmetrical middle range $|\omega|<V/2$. Since the resonant frequencies of the transmission channels through the finite middle chain are in symmetrical pairs ±$\pm\omega_{n}$, raising $V$ effectively only opens new channels in pairs, producing only $M+1$ conductance steps. In the quantum gas scenario with particles coming only from the left reservoir, in contrast, raising $\mu_{L}$ through the system’s frequency band sweeps through each of the $12M+$ 1 transmission resonances separately.
Figure 3: Dimensionless fermion current $I$ at zero temperature, as a function of left-side chemical potential $\mu_{L}$, for two different central-site loss rates $\gamma$. Both plots show $\tau_{1}=0.1$ as in Fig. 2, but with $V=0$, $\Lambda=1$, and no fermions incident from the right. Although the middle chains here have seven sites ($M=3$), the leftmost and rightmost current steps are small and gentle in a), and completely flattened in b), leaving only five current steps clearly visible. In a) we have $\gamma=0.5$, while b) shows $\gamma=5$. As well as suppressing the outermost of the seven current steps, the moderate loss in a) has noticeably smoothed out two of the five more visible steps in the current, while the stronger loss in b) has smoothed them completely into gradual slopes. The other three visible current steps are not significantly affected by loss in either case, since they are due to transport through odd-parity channels having nodes at the lossy site $m=0$. Although $I(\mu_{L})$ differs in detail from analogous plots of $I(V)$ for quantum dots between leads, the same phenomena of conductance steps with parity-dependent smoothing by loss are revealed.
@@
Five steplike rises in conductance through the finite middle chain of seven sites ($M=3$) are clearly visible in Fig. 3; looking closely at a), one can also discern two more small, smooth steps at the left and right edges of the plot. Also apparent is the effect of loss in making some steps noticeably steeper than others. In Fig. 3a) the tiny first and seventh steps, as well as the more visible third and fifth steps, are all due to even-parity channels that are affected by the loss at $m=0$, and so more gently sloped than the three other steeper steps, which are due to odd-parity channels with nodes at $m=0$. The ten-times-higher loss rate in Fig. 3b) is enough to suppress the first and seventh steps entirely, and eliminate the shoulders of what were the third and fifth steps in a), leaving them as gradual slopes, while still leaving the three parity-protected transport resonances as sharp steps.
@@ IV.2 Finite temperature effects
Since the main phenomena to be studied in this kind of transport are the conductance steps and their parity- and $\gamma$-dependent steepness, thermal smearing of the conductance curve is a serious issue for experiments. Since the frequency spacing of our transmission resonances is of order $2/(2M+1)$, temperatures higher than this spacing (in units of $\hbar\tau/k_{B}$) will wash out the conductance steps. Even if $T/\tau$ is below the $2/(2M+1)$ threshold, moreover, thermal smoothing of the conductance steps can reduce their dramatic signature of discrete transmission channels to only a slight waviness in $I(\mu_{L})$, as seen in Fig. 4a); in this regime, moreover, the additional smoothing of even-parity channels due to the $m=0$ loss can be hard to discern. Both the steps and the effect of parity can still be seen at sufficiently low temperatures, as in Fig. 4b), but such low temperatures may be experimentally challenging. If the basic $T<2\tau/(2M+1)$ limit can be beaten, the problem of losing visibility of the parity effect for $T\lesssim\gamma$ can be solved by increasing $\gamma$, as in Fig. 4c).
@@ V.2 Quasi-stationary currents
For example, a natural initial state to prepare experimentally is a finite-system analog of the current-from-the-left scenarios shown for infinite lattices in Fig. 3. We can prepare the initial state in the same way that we described above for the infinite system, by initially setting $\tau_{1}\to 0$ to fully break the weak link and isolate the middle chain of our lattice from its outer ends. We then populate the isolated left end of the lattice with fermions in equilibrium, leaving the middle and right-end lattices empty. Since adiabatically ramping up $\tau_{1}$ would still just lose all of our finite number of particles, however, we instead abruptly restore non-zero $\tau_{1}$ $t=0$, and let the left-end population of fermions flow through the weakly linked middle chain, with its loss at $m=0$, and into the right lattice end. At different finite times $t>0$ we measure the decreasing population difference between left and right sides,
@@
As we can see in Fig. 5), there is a brief initial period in which the fermions first begin to move through the middle chain of sites, via the weak links. Within a time that is at most of order $M/\tau$, however, this initial onset phase ends and we can see a remarkably steady fall of $\Delta\mathcal{N}$. The initial equilibrium state in the left end of the lattice can be viewed as an ensemble of wave packets moving in both directions; these packets move to and through the middle chain from further back in the left reservoir, just as they would if the reservoir were infinite. Even the packets which begin moving to the left, and then approach the middle chain after reflecting from the left end of the lattice, simply reproduce the effect of incoming packets from beyond the lattice end. The first effect which does not look just like an infinite system, in fact, is the return to the weak link of packets which have been reflected from it, and then reflected again from the left end of the lattice. This occurs for packets with group velocity $v_{G}$ at time $t=2N/v_{G}$. The maximum group velocity in our system is $2\tau$, occurring at $\omega=0$, and so the earliest finite size effects are those that appear in Fig. 5 at $t=N/\tau$. Between $t\gtrsim M/\tau$ and $t<N/\tau$ there can be plenty of time to observe a quasi-stationary current, as long as $N$ is sufficiently larger than $M$. And as Fig. 5 shows, the rate at which $\Delta\mathcal{N}$ falls, during this quasi-stationary epoch between the initial transient and the onset of finite-size effects, can closely match the steady-state current of the infinite system.
@@
Figure 5: Finite-time evolution in finite systems. The initial state is prepared in a finite lattice with $N=20$, 40, or 80 sites in the outer lattice on each side. With the weak links between reservoir ends and middle chain initially turned off completely ($\tau_{1}\to 0$), the $N$-site left end lattice is initially half-filled at zero temperature, with the right end and central chain of seven sites ($M=3$) both initially empty. At $t=0$ the weak links are turned on $\tau_{1}\to 0.1$ and the system’s single-particle density matrix is evolved in time under (II.3.2) numerically, for a duration $2.5N/\tau$. The difference in total left- and right-side populations $\Delta N/2=(N_{L}-N_{R})/2$ is plotted in each case of $N$. The dashed line has the slope $-I$, where $I$ is the steady-state current for the $N\to\infty$ limit as shown in Fig. 3. The early transient epoch is so brief that it can only barely be seen with the shorter time axis of the $N=20$ plot. The onset of finite-size effects at $t=N/\tau$ is clear in all cases: with group velocity $2\tau$ for half filling, this is how long it takes for particles that initially reflect from the weak link to return to the weak link after reflecting again from the end of the lattice. For times after the short initial transient and before $t=N/\tau$, the finite system’s particle difference closely reproduces the infinite system’s steady-state current—in the case $\mu_{L}=0$ of initial half-filling. See Fig. 6 for other cases of $\mu_{L}$.